OpenStudy (anonymous):

2 sandwiches and 1 drink is $7.50. 5 sandwiches and 2 drink is $18.00. How much is each sandwich? Write the correct equations for this word problem and points for solving it correctly and answering the the question in a complete sentence

OpenStudy (anonymous):

The easiest way and really the only way to solve this is by using equations . Let sandwhiches be represented by x and drinks by 7. 2x + y = 7.50 5x + 2y= 18.00. Now what you need to do is get one variable....in this case we are going to use just x. 2x + y = 7.50. Subtract 2x on both sides and you know y = 2x -7.50 . So plug that in to the other equation and you have 5x +2(2x-7.50)= 18.00 and now you just solve that.